Let c,r be constants, and D={(x,y,z):x^2+y^2+z^2≤r^2}. The answer to ∭D cdV

is
Select one:
a. (π^2cr^3)/3


b. (4πcr^3)/3

c. 4πcr^3

d. (πcr^4)/3


e. (4πcr^2)/2

f. (4r^3)/3


g. (πcr^3)/3
1
Expert's answer
2020-06-03T19:24:31-0400

Correct option is (b).

Reason:



Given ,"c,r" are constants.

Assume radius is "R" instead of "r" for the time being,then replace "R" by "r" in the final result.

"D" is the region described as "D=\{(x,y,z):x^2+y^2+z^2≤r^2\}" ,clearly "D" contains all the points lies inside and surface of the sphere whose radius is "r" .

Thus,

"I=\int \int\int_{D}cdV=c\int \int\int_{D}dV"

Let's draw the elemental volume


Hence,

"I=c\int \int\int_{D}dV=c\int_{0}^{R}\int_{0}^{\pi}\int_{0}^{2\pi}r^2\sin(\theta)d\theta d\phi dr\\
\implies I=c\int_{0}^{R}\bigg(r^2\int_{0}^{\pi}\bigg(\sin(\theta)\int_{0}^{2\pi}d\phi \bigg)d\theta \bigg)dr=c\frac{4\pi R^3}{3}"

Therefore the final answer is

"I=c\frac{4\pi r^3}{3}"